OpenStudy (anonymous):

Solve the equation –8 = 3(–x + 5) – x and check

OpenStudy (anonymous):

–8 = 3(–x + 5) – x -8 = -3x+15-x -8 = -4x+15 -8-15 = -4x -23 = -4x x = 23/4 just plug back into the original equation to see if it is a solution.